# Data Management in the Noisy Intermediate-Scale Quantum Era

### v1

@@ 4.2. Databases to the rescue
At its core, a CQSS must, at a minimum, be capable of evaluating quantum circuits. This primarily involves performing linear algebraic operations, often described in terms of tensor networks. [^3] For simplicity of exposition, we do not consider non-tensor-based simulation methods which represent quantum states by their symmetries, rather than by complex-valued vectors (Gottesman, 1998). A tensor is a multidimensional array, and the dimensions along which a tensor extends are its indices. A tensor network is a collection of tensors represented as nodes, with the shared indices contracted. Tensor contraction is the summation of shared indices between tensors and is a generalization of matrix multiplication (Biamonte, 2020), as shown in Fig. 2c. The first question that comes to mind is as follows.
@@
In other words, should we map the simulation workload to SQL queries and store the results in a DBMS? The advantage is that we can benefit from the efficiency and portability of modern database systems. And indeed, initial efforts from the database community (Blacher et al., 2023; Trummer, 2024) have begun to address the out-of-core simulation challenge by mapping quantum states and gates to SQL queries. For instance, in (Blacher et al., 2023), qubit states and gates are represented as tensor networks, with sparse tensors in COO format [^4] https://docs.scipy.org/doc/scipy/reference/generated/scipy.sparse.coo_matrix.html and Einstein summation operations mapped to SQL queries. Additionally, the recent abstract (Trummer, 2024) introduces the idea of supporting dense vectors using SQL UNION ALL and CASE statements. However, despite these early efforts that support basic quantum states and gates up to 18 qubits and a circuit depth of 18 in separate experiments, the core challenges of representing general quantum states and operations in DBMSs and achieving scalability remain unresolved. Furthermore, preliminary findings question whether query optimizers in existing RDBMSs can effectively optimize SQL queries for simulation workloads (Blacher et al., 2023). Therefore, designing intermediate representations (Karanasos et al., 2020; He et al., 2022) and seeking more effective query optimization, potentially beyond existing techniques (Neumann, 2024), could be crucial and require further exploration.
On the other hand, the connection to tensor computations suggests the following question.
@@
While both machine learning and quantum applications share fundamental building blocks like matrices, tensors, and tables, we must address challenges unique to quantum computing. We highlight two key gaps that need to be overcome:
@@ 4.2.1. Quantum-specific optimizations
Existing simulators for quantum computing are mostly implemented in Python (cuQuantum team, 2024; Vincent et al., 2022), C++ (AI, 2021; Vincent et al., 2022) or Julia (Luo et al., 2020). There are also dedicated quantum programming languages and domain-specific compilation techniques. See the survey (Heim et al., 2020). In addition, tensor network rewriting techniques have been considered in the ZX-calculus (Coecke and Duncan, 2011). For tensor-based simulators, a defining feature is the optimization of contraction order, finding the optimal contraction path (Young et al., 2023; cuQuantum team, 2024). We refer to (Young et al., 2023) for a more in-depth exposition of simulators and their optimizations. We will need to incorporate the simulation-specific optimizations in our CQSS as well, and in such a way that these also operate out-of-core. This may incur interesting algorithmic challenges. The following question now arises.
@@
Such an optimizer should, given a circuit, determine the most efficient way to simulate it. For example, when a circuit allows for a tractable simulation (e.g., circuits of bounded treewidth), the optimizer should compile it into an efficient simulation. Similarly, based on the sparsity of (intermediate) quantum states, specialized sparse gate computations may be employed instead of the standard dense computations. Particular to the quantum setting is the noisy character of the quantum computation, which is known to impact the efficiency of simulation (Aharonov et al., 2023; Qi et al., 2020; Pan and Zhang, 2021; Huang et al., 2020, 2021). Hence, an optimizer needs to take noise levels into account. Moreover, the optimizer should consider the downstream task – whether it’s computing precise probabilities for strong simulations or sampling measurement outcomes for weak simulations – and select the appropriate simulation strategy (algorithm) accordingly.
@@ 4.2.2. Quantum-specific data representations
An operation on a data structure is tractable if it executes in polynomial time with respect to the input size (Darwiche and Marquis, 2002). The choice of data representation can significantly impact whether certain quantum operations (e.g., gates, measurement explained in Sec. 2.2) are tractable (Vinkhuijzen et al., 2024). We emphasize three main requirements for data representation: expressiveness, closure, and succinctness/tractability.
For expressiveness, the data representation must be sufficiently rich to represent quantum states and measurements. Closure refers to the property that the result of quantum operations on the represented states can also be represented within the same data representation. This property ensures compositionality. Finally, it is crucial to minimize the space required to represent a quantum state, as the number of qubits and the level of entanglement can result in exponentially large states. Therefore, we seek data representations that are succinct, minimizing the space needed to store data while still allowing for efficient query processing. It may also be interesting to consider these questions beyond tensor network representations as well (Bravyi and Gosset, 2016; van de Wetering, 2020; Zulehner and Wille, 2018; Jozsa and Miyake, 2008).
By addressing these challenges, we can create a robust and efficient classical-quantum simulation system that integrates the strengths of existing optimization techniques of databases, tensor computation, and simulations, while meeting the specific needs of quantum computing. Due to space limitations, we will not detail the consistency and recovery aspects of the envisaged CQSS; We comment on its scientific data management part in the next section.
@@ 4.3. More data management opportunities
Quantum error correction & Graph analytics. As discussed in Sec. 2.1, qubits in the NISQ era are inherently vulnerable to noise, which can significantly impact the reliability of computation results. Quantum error correction is crucial for ensuring reliable quantum computation. This process consists of two main phases: coding and decoding. In the coding phase, the input quantum information is coded into a quantum error-correcting code such as the surface code (Fowler et al., 2012; Bravyi et al., 2018; Fowler et al., 2009). During the decoding phase, errors are detected and corrected. Typically, errors are modeled as graphs. The error detection and correction algorithms, referred to as decoders, are often built upon graph theory. For instance, in certain promising error-correction schemes (Fowler, 2015; deMarti iOlius et al., 2024; Wu et al., 2022), decoding involves solving a minimum-weight perfect matching (MWPM) problem on a graph where each node represents a check measurement. Exploring the recent effort of graph analytics (Manolescu and Mohanty, 2023; Yan et al., 2005; Wang et al., 2021; McGregor, 2014; Behnezhad et al., 2023) and graph databases (Tian, 2023) such as Neo4j [^5] https://neo4j.com/, TigerGraph [^6] https://www.tigergraph.com/, might yield valuable insights to advance these error correction techniques. A unique challenge in graph data management for error correction is balancing accuracy and efficiency (deMarti iOlius et al., 2024). A more accurate decoder generally requires higher computational complexity and, therefore, operates more slowly. However, the fidelity of a quantum state decays over time, demanding fast error detection and correction by the decoder.
Quantum experiments & Scientific data management. Quantum experiments, like any other scientific experiments, generate data and require data management. For instance, quantum mechanics experiments on supercomputers often apply HDF5 files to store data and XML files for metadata (Cruz et al., 2012). Typical challenges include avoiding I/O bottleneck (efficient I/O parallelization) and handling the sheer volume and heterogeneity of data files. The scientific data management challenge is also domain-dependent, such as those arising from quantum chemistry or condensed matter physics experiments. These challenges may require us to push forward the frontier of existing database topics such as metadata management (Bernstein et al., 2000; Amano et al., 2014) and data formats (Zeng et al., 2023). Moreover, it is also interesting to explore how to build novel data lakes (Nargesian et al., 2019; Hai et al., 2023) or lakehouses (Armbrust et al., 2020, 2021; Jain et al., 2023) for quantum applications. All these challenges also relate to enabling and improving the simulation workflow of our envisioned classical-quantum simulation system, as described before.
Quantum network & Distributed databases. Quantum networks hold the potential to offer advantages and capabilities beyond those of classical internet systems, e.g., secure communication (Broadbent et al., 2009) or distributed computing (Nickerson et al., 2014). Developing quantum network requires simulators of various kinds, such as simulating timing behaviour accurately (Coopmans et al., 2021; Wu et al., 2021) or for supporting the development of application software (Dahlberg and Wehner, 2018). It is interesting to explore whether the quantum network applications and tools can be supported and advanced by distributed databases (Özsu et al., 1999) and cloud-native databases (Dong et al., 2024), while also considering quantum-specific features like entanglement.

### v2

@@ 4.2. Databases to the rescue
At its core, a CQSS must, at a minimum, be capable of evaluating quantum circuits. This primarily involves performing linear algebraic operations, often described in terms of tensor networks (Biamonte, 2019). [^3] For simplicity of exposition, we do not consider non-tensor-based simulation methods which represent quantum states by their symmetries, rather than by complex-valued vectors (Gottesman, 1998). A tensor is a multidimensional array, and the dimensions along which a tensor extends are its indices. A tensor network consists of vertices representing tensors, and edges the indices. Free indices are depicted as legs, i.e., edges that only connect to one vertex in the graph. Connecting two vertices by joining a leg corresponds to the contraction with the corresponding indices. Fig. 2b shows an example of the tensor network representation of GHZ state. Tensor contraction is the summation of shared indices between tensors and is a generalization of matrix multiplication (Biamonte, 2020). See Appendix A of (Hai et al., 2025) for more details. The first question we ask is as follows.
[…]
(a) GHZ State (sparse circuit)
(b) QFT (dense circuit)
Figure 3. Memory usage comparison of three RDBMS-based solutions (red) against a general approach, i.e., order-n tensor baseline (blue) and the SotA MPS baseline (green). The x-axis shows increasing numbers of qubits, and the y-axis shows memory consumption (KB) in $\log_{10}$-scale. For sparse circuit (GHZ), RDBMS solutions demonstrate significantly lower memory usage than the order-n tensor baseline, with SQLite showing a slight advantage over the MPS baseline. For the dense circuit QFT, the memory usage of RDBMS solutions grows comparably to or exceeds the order-n tensor baseline, indicating the need for improvements in handling dense tensor computations.
@@
In other words, should we map the simulation workload to SQL queries and store the results in a DBMS? The advantage is that we can benefit from the efficiency and portability of modern database systems. And indeed, initial efforts from the database community (Blacher et al., 2023; Trummer, 2024) have begun to address the out-of-core simulation challenge by mapping quantum states and gates to SQL queries. For instance, in (Blacher et al., 2023), qubit states and gates are represented as tensor networks, with sparse tensors in COO format [^4] https://docs.scipy.org/doc/scipy/reference/generated/scipy.sparse.coo_matrix.html and Einstein summation operations mapped to SQL queries. Additionally, the recent abstract (Trummer, 2024) introduces the idea of supporting dense vectors using SQL UNION ALL and CASE statements. Despite these efforts that support basic quantum states and gates up to 18 qubits and a circuit depth of 18 in separate experiments, the core challenges of representing general quantum states and operations in DBMSs and achieving scalability remain. Moreover, our preliminary results in Fig. 3 demonstrate that RDBMS-based simulations are efficient for quantum circuits involving sparse tensor computations. All experimental details, together with a space and time complexity analysis, can be found in Appendix B of (Hai et al., 2025). RDBMS-based solutions are competitive with, and may even outperform, state-of-the-art representations like matrix product states (MPS) (Pérez-Garcia et al., 2007) in specific scenarios. Interestingly, this advantage does not extend to dense circuits such as quantum fourier transform (QFT) circuits, highlighting the need for further optimizations to enhance RDBMS systems for simulation workloads. In Appendix C of (Hai et al., 2025) we additionally discuss the use of NoSQL databases. On the other hand, the connection to tensor computations suggests the following question.
@@
To address challenges unique to quantum computing, we highlight two key gaps that need to be overcome:
@@ 4.2.1. Quantum-specific optimizations
Existing simulators for quantum computing are mostly in Python (cuQuantum team, 2024; Vincent et al., 2022), C++ (AI, 2021; Vincent et al., 2022) or Julia (Luo et al., 2020). There are also dedicated quantum programming languages and domain-specific compilation techniques. See the survey (Heim et al., 2020). In addition, tensor network rewriting techniques have been considered in the ZX-calculus (Coecke and Duncan, 2011). ZX-calculus is a graphical language for representing and reasoning about quantum computations, represented as tensor networks with specialized tensors called spiders (van de Wetering, 2020b). Its graphical nature enables simplifications by locally merging connected vertices while preserving the original tensor, useful in various contexts including formal verification (van de Wetering, 2020b; Lehmann et al., 2022; de Beaudrap and Horsman, 2020).
Quantum-specific optimizations share some conceptual similarities with classical database optimization, yet they diverge significantly due to the unique characteristics of quantum data and operations. First, classical databases rely on relational or NoSQL data models (graph, document-based, etc.) with indexing mechanisms like B-trees or hash tables to facilitate efficient access and retrieval. In simulations, however, quantum states are represented as state vectors (Fig. 2a) or by using more advanced data structures such as tensor networks (Fig. 2b). Second, in classical databases, query optimization involves techniques like join ordering to identify the optimal ordering of join operations between relations for an efficient query plan (Steinbrunn et al., 1997; Leis et al., 2015; Trummer and Koch, 2017). For tensor-based simulators, a defining feature is the optimization of contraction order, which determines the most efficient sequence to contract tensors representing quantum states (Young et al., 2023; cuQuantum team, 2024). Contraction order optimization parallels the role of join order optimization in traditional databases but is more complex due to the exponential growth of quantum state dimensions and the unique features of quantum data, such as entanglement. Third, classical database systems rely heavily on cost estimation to predict and minimize resource use for query execution. Quantum computation simulators, however, lack cost estimation frameworks and instead use optimizations like parallelization, SIMD (Single-Instruction, Multiple-Data) processing, and matrix decomposition techniques (e.g., SVD) to improve performance (Young et al., 2023). Quantum-specific issues, such as quantum noise (Sec. 2), complicate simulation further. These differences lead to the need for fundamentally new optimization frameworks to manage and simulate quantum computation effectively, beyond classical structures and processes. The following question now arises.
@@
An optimizer can be designed, given a circuit, to determine the most efficient way to simulate it. For example, when a circuit allows for a tractable simulation (e.g., circuits of bounded treewidth), the optimizer should compile it into an efficient simulation. Similarly, based on the sparsity of (intermediate) quantum states, specialized sparse gate computations may be employed instead of the standard dense computations. Particular to the quantum setting is the noisy character of the quantum computation, which is known to impact the efficiency of simulation (Aharonov et al., 2023; Qi et al., 2020; Pan and Zhang, 2021; Huang et al., 2020, 2021). Hence, an optimizer needs to take noise levels into account. Moreover, the optimizer should consider the downstream task – whether it’s computing precise probabilities for strong simulations or sampling measurement outcomes for weak simulations – and select the appropriate simulation strategy (algorithm) accordingly.
@@ 4.2.2. Quantum-specific data representations
An operation on a data structure is tractable if it executes in polynomial time with respect to the input size (Darwiche and Marquis, 2002). The choice of data representation can significantly impact whether certain quantum operations (e.g., gates, measurement explained in Sec. 2 and Appendix A of our technical report (Hai et al., 2025)) are tractable (Vinkhuijzen et al., 2024). We emphasize three main requirements for data representation: expressiveness, closure, and succinctness/tractability.
For expressiveness, the data representation must be sufficiently rich to represent quantum states and measurements. Closure refers to the property that the result of quantum operations on the represented states can also be represented within the same data representation. This property ensures compositionality. Finally, it is crucial to minimize the space required to represent a quantum state, as the number of qubits and the level of entanglement can result in exponentially large states. Therefore, we seek data representations that are succinct, minimizing the space needed to store data while still allowing for efficient query processing. It may also be interesting to consider these questions beyond tensor network representations as well (Bravyi and Gosset, 2016; van de Wetering, 2020a; Zulehner and Wille, 2018; Jozsa and Miyake, 2008). A concrete idea for a novel data representation is to combine MPS and the recently-proposed local-invertible map decision diagram (LIMDD) (Vinkhuijzen et al., 2023). LIMDD compresses a state vector by lumping together parts of the vector that are equivalent modulo simple quantum gates. In polynomial time and space in the number of qubits, LIMDDs can simulate circuits that MPS cannot and vice versa (Vinkhuijzen et al., 2024). One approach to combine these strengths in a single data structure is applying the lumping to the MPS matrices instead of quantum state vectors, by building upon existing work on equivalence characterizations of MPS (Lami and Collura, 2024).
{myframe}
[roundcorner=3pt, innerleftmargin = 5pt] Q5. Are there other benefits from relying on database technology?
By using database technology we aim for the simulation of complex circuits that require significant memory, without relying on HPC infrastructures or the operating system’s memory management. By controlling what data is pushed to secondary storage during simulations, we can achieve more efficient I/O behavior. Additionally, transaction management and recovery become important—where a transaction could, for example, represent a sequence of quantum gates—ensuring that computations can restart from saved checkpoints while maintaining the correctness of partially saved results. Parallel evaluation strategies in quantum circuit simulations also highlight the need for effective transaction scheduling, leading to more reliable computational processes. Undoubtedly, many challenges remain in this context when considering quantum simulation. We prioritize Q1–Q4, however, as first steps. A more extensive discussion can be found in Appendix D in (Hai et al., 2025).
@@ 4.3. More data management opportunities
Quantum error correction & Graph analytics. Quantum error correction (QEC) enables reliable execution of quantum computation on noisy quantum processors and is a crucial part of the roadmap to fault-tolerant quantum computation (Terhal, 2015). QEC process consists of two main aspects: coding and decoding. The input quantum information as well as the operations to be performed are coded using a quantum error-correcting code such as the surface code (Fowler et al., 2012; Bravyi et al., 2018; Fowler et al., 2009). Intermittently, decoding is applied: errors are detected and corrected. Decoders  (Fowler, 2015; deMarti iOlius et al., 2024; Wu et al., 2022) often leverage graph theory, solving tasks like a minimum-weight perfect matching (MWPM) problem on a graph where each node represents a check measurement. For MWPM, the number of nodes is of the same order as the number of qubits. Many interesting research problems arise how to model such graphs and design the data formats. Moreover, QEC has to be performed at high speeds to avoid decaying quantum states over time (see Sec. 2), e.g. at most several microseconds for some types of quantum hardware for decoding (Battistel et al., 2023). The scalability to many qubits, and speed required for QEC might benefit from advanced data management tools and techniques (Manolescu and Mohanty, 2023; Yan et al., 2005; Wang et al., 2021; McGregor, 2014; Behnezhad et al., 2023), offering a promising direction for further exploration.
Quantum experiments & Scientific data management. Quantum experiments, like any other scientific experiments, generate data and require data management. For instance, quantum mechanics experiments on supercomputers often apply HDF5 files to store data (Cruz et al., 2012). Another compelling direction is to explore how to build specialized data lakes (Nargesian et al., 2019; Hai et al., 2023) or lakehouses (Armbrust et al., 2020, 2021; Jain et al., 2023) to support scientific data management for quantum computing.
ML & Quantum data management. i) Simulation: as explained in Sec. 4.2.1, a key challenge in optimizing tensor network based simulation is to find optimal tensor contraction order. A recent research direction is to apply machine learning (ML), specifically reinforcement learning (RL) and graph neural networks (GNN), to optimize tensor contraction order, addressing the computationally intensive nature of this challenge (Meirom et al., 2022; Liu and Zhang, 2023). ii) Error correction: decoding methods for quantum error correction (QEC), like MWPM, face scalability challenges in large quantum systems, where rapid error detection within strict time limits is essential (Vittal et al., 2023). An interesting new direction is data-driven QEC, which employs ML techniques to quantum error correction, such as RL (Zeng et al., 2023; Nautrup et al., 2019; Andreasson et al., 2019), multilayer perceptrons (MLP) (Chu et al., 2022), convolutional neural networks (Chamberland et al., 2023), and GNN (Lange et al., 2023). iii) Improving quantum algorithm efficiency: Quantum algorithms are represented and implemented as quantum circuits, where efficiency can be improved by reducing costly gates like SWAP gate or by minimizing circuit depth and gate count. ML methods, specifically RL (Fan et al., 2022; Fösel et al., 2021) and MLP (Paler et al., 2023), have shown potential in optimizing circuit design to enhance the efficiency of quantum circuits on real devices.